(6) Determine the truth value of the statement ∃x∀y(x lessthanorequalto y2) if the domain for the
variables consists of
(a) The positive real numbers.
(b) The integers.
(c) The nonzero real numbers.
1
Expert's answer
2020-09-14T16:52:42-0400

We have the predicate "P(x,y)="x\leq y^2"" (two variables). We have (a) "\R^+" as a domain for both of the variables. We should determine for which values of "x" and "y" the predicate "\exist x\forall y (x\leq y^2)=1."

Consider the predicate "S(x)=\forall y (x\leq y^2)."

Here "y" should be greater or equal to "\sqrt{x}". "\text{Then } S(x)=1, \exist x S (x)=1.\\"

So for all positive real numbers "x" and "y\geq \sqrt{x}" the predicate "\exist x\forall y (x\leq y^2)=1."

(b) "\Z"

If "x\geq 0" then "y\geq \sqrt{x}."

If "x<0" then "\forall y."

(c) "\R-\{0\}"

If "x>0" then "y\geq \sqrt{x}."

If "x<0" then "\forall y."